# SL2A-INR: Single-Layer Learnable Activation for Implicit Neural Representation

### v3

@@ Abstract
Implicit Neural Representation (INR), leveraging a neural network to transform coordinate input into corresponding attributes, has recently driven significant advances in several vision-related domains. However, the performance of INR is heavily influenced by the choice of the nonlinear activation function used in its multilayer perceptron (MLP) architecture. To date, multiple nonlinearities have been investigated, but current INRs still face limitations in capturing high-frequency components and diverse signal types. We show that these challenges can be alleviated by introducing a novel approach in INR architecture. Specifically, we propose SL2A-INR, a hybrid network that combines a single-layer learnable activation function with an MLP that uses traditional ReLU activations. Our method performs superior across diverse tasks, including image representation, 3D shape reconstruction, and novel view synthesis. Through comprehensive experiments, SL2A-INR sets new benchmarks in accuracy, quality, and robustness for INR. Our Code is publicly available on GitHub.
@@ 1 Introduction
Implicit neural representations (INRs), which model continuous functions over discrete data points, have recently garnered significant attention for their ability to effectively represent 2D images, 3D shapes, neural radiance fields, and other complex structures [23, 34, 26, 32]. Unlike traditional discrete grid-based methods, INRs focus on training neural networks, particularly Multilayer Perceptrons (MLPs), that utilize continuous element-wise nonlinear activation functions to map continuous input coordinates to their respective output values, thereby delivering a continuous signal representation [34]. Due to their flexible, compact, and efficient design, INRs have garnered immense interest in solving tasks dealing with complex and high-dimensional data. They have demonstrated versatility across a wide range of applications, including computer graphics [23, 24, 6], computer vision [20, 45], inverse problems [38], and signal representations [42, 34]. In contrast to Convolutional Neural Networks (CNNs), INRs are not constrained by locality biases, thus exhibit exceptional generalization abilities, feature a concise formulation, and display enhanced flexibility. However, their potential is constrained by various obstacles that limit the application of INRs. Conventional neural networks such as ReLU-based MLPs struggle to capture high-frequency details accurately in signals, such as intricate textures in images or complex shapes in 3D reconstruction, due to their inherent bias towards learning simpler, lower-frequency patterns, known as spectral or low-frequency bias [27, 43, 14]. The pioneering efforts in mitigating this issue can be roughly grouped into several main categories. One approach involves increasing the complexity of the input data manifold and explicitly extracting high-frequency features, such as by positional encoding [39, 24]. An alternative major category of methods utilizes complicated activation functions to achieve more precise approximations [34, 10, 31]. Despite promising results, these approaches still exhibit some degree of spectral bias [44], primarily learning faster lower-frequency components of the input. This indicates that there is still room for improvement in addressing spectral bias more comprehensively [5]. Moreover, periodic activation functions are particularly vulnerable to initialization methods, experiencing notable performance declines when not strictly following initialization protocols [29]. Therefore, these approaches have difficulty representing fine details in complex shapes, limiting their flexibility across signal types and inverse problems. To overcome these challenges, we propose a novel approach that significantly boosts the hierarchical representation capabilities of INRs. Our method achieves superior performance in high-fidelity reconstructions across a wide array of tasks, including images and intricate 3D structures. Additionally, it addresses complex problems such as Neural Radiance Fields (NeRF).
@@ 2 Related Works
Implicit Neural Representation Recent studies have demonstrated significant success in utilizing neural networks to represent various types of signals. These representations are applied across a range of applications, including images [29, 31, 34], occupancy volumes [25], view synthesis [23, 2, 11], and audio/video signals [37, 7]. However, the widely-used ReLU activation function has limitations in learning low-frequency information, which can affect the performance of standard MLPs [39]. In this context, the ReLU activation function often fails to accurately represent the signal, resulting in the loss of important high-frequency components and leading to a suboptimal overall representation. To address this issue, several solutions have been proposed in the literature. One approach involves encoding the input data with sinusoids [39], which maps the input coordinates to a higher dimensional space. The output of this positional encoding is then used as input to the MLP. Another line of research focuses on modifying the activation functions of the MLP. For instance, using periodic activations, as in SIREN [34], Gaussian activations [29], or wavelet-based activations [31], or variable-periodic activation functions [19] can help capture high-frequency details more effectively. However, these methods also come with their own challenges. For example, the representational capacity of SIREN can be highly dependent on the choice of hyperparameters for sinusoidal functions, such as frequency, and can be sensitive to initialization, necessitating careful design choices to avoid random variations [41].
@@
Learnable Activation Functions As previously noted, the methods for mitigating spectral bias in INRs have introduced new challenges. In particular, networks utilizing custom activation functions, such as periodic activations, are highly sensitive to initialization schemes and hyperparameter choices. Specifically, the inherent periodicity of sine functions can cause oscillations and slow down convergence, creating challenges during the training process [19]. Additionally, despite advancements, some degree of spectral bias remains present in these approaches. Consequently, it is critical to develop an adaptive function to handle nonlinearity and complex frequency distribution effectively. Instead of relying solely on custom-designed activation functions, which may have their own limitations, we propose using learnable activation functions that adapt based on the data itself as part of our network. A learnable activation function is a type of activation function in neural networks that is not fixed but is optimized during training. Unlike traditional activation functions, which are predetermined and static, learnable activation functions have parameters that are adjusted through backpropagation [13, 3]. These functions can be represented by various forms, such as polynomials, splines, or sigmoid linear units [13, 9, 28], and are designed to adapt dynamically to the data and learning process, potentially enhancing the network’s ability to capture complex patterns and improve performance. Specifically, the recent Kolmogorov-Arnold Network (KAN) [18] utilizes B-splines for their activation functions. Capitalizing on this line of research, we transform the input coordinates into a high-dimensional space by projecting them onto multiple adaptive Chebyshev polynomial [36] functions of a fixed degree, departing from the fixed activation functions or positional encodings. These polynomials act as learnable activations, allowing the network to adaptively capture a wide spectrum of frequency information by leveraging the spectral properties of Chebyshev polynomials to efficiently represent both low and high-frequency components of the input signal (as they exhibit progressively higher-frequency oscillatory components in higher degrees). We provide intuitive justifications for Chebyshev polynomial design choices throughout the paper. Besides, we incorporate a feature interaction mechanism where the Learnable Activation (LA) Block’s output serves as a dynamic modulator for subsequent layers, enhancing the model’s ability to capture and propagate high-frequency features throughout the network. For simplicity, we use SL2A and SL2A-INR interchangeably throughout the paper to refer to our model.
@@ 3.2.2 Fusion Block
Table 1: Comparison of different methods based on PSNR (in dB) / SSIM (0-1) for 2D color image approximation results for images 00 to 07. Higher values indicate better quality, with red indicating the best and blue indicating the second best.
Method #Params (K)$\downarrow$ 00 01 02 03 04 05 06 07
@@
Table 2: Continuation of Tab. 1. Comparison of different methods based on PSNR (in dB) / SSIM (0-1) for 2D color image approximation results for images 08 to 15. Higher values indicate better quality, with red indicating the best and blue indicating the second best.
@@
where $g(x)$ is defined over 2048 discrete points, uniformly sampled from the interval $[-1,1]$. In this expression, $R$ represents the rounding function, which introduces discontinuities into the data and increases the difficulty of training. This setup enables us to analyze how the network handles different frequency components, particularly in the presence of spectral bias.
@@ 5 Experimental Results
To evaluate the effectiveness of our proposed method, we conducted several experiments in signal representation tasks such as 2D image representation, 3D shape reconstruction, and novel view synthesis (presented in supplementary materials). Our experiments were conducted using PyTorch on an Nvidia RTX 4070 GPU equipped with 12GB of memory. We explored all possible combinations of learning rates (1.5$\times$10-4, 10-4, 1.5$\times$10-3, 10-3, 10-2) and batch sizes (32$\times$32, 64$\times$64, 128$\times$128, 256$\times$256). For each combination, we ran each model on five different images [40], averaging the results to select the best configuration of learning rate and batch size.
@@ 5.1.1 Image Representation
We first applied our method to the task of 2D image fitting using INR. In our model, the first layer employs Chebyshev polynomials ($K$=256), while the subsequent layers use simple ReLU activation functions. The loss function used is the MSE between the model’s output and the actual pixel values in the images.
@@
Results We evaluated our method by comparing its performance against six baseline models in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [15]. The baselines included WIRE [31], SIREN [34], FINER [19], ReLU+P.E. [39], and Gauss [29]. For consistency, we trained each model for 500 epochs and reported the best results achieved. All models were implemented with a standard architecture of $L=5$ layers, each containing 256 neurons, except for WIRE, which, following its original design, used a multilayer perceptron (MLP) with 4 layers and 300 neurons per layer. In our approach, for a comparable number of parameters, we used a Learnable Activation (LA) Block with polynomial degrees set to $K=256$, and incorporated low-rank linear layers with rank 128 in the Fusion Block followed by ReLU activations. This architecture was designed to balance flexibility and computational efficiency, enabling the model to effectively capture high-frequency details while maintaining efficient computation.
The quantitative results for images 00 to 07 are shown in Tab. 1, while the results for images 08 to 15 are presented in Tab. 2. Our model performed best, or at least second-best, in most images, achieving state-of-the-art results. In the few instances where the results were comparable, our approach still demonstrated its effectiveness, achieving competitive performance. Furthermore, our model produced the best average results across all images, highlighting its overall effectiveness. We also provide qualitative results in Fig. 3, which includes all images with their corresponding baseline methods, as well as our model. The figure contains magnified regions of the images where text is present. The zoomed-in portions show that our approach outperforms the baselines by reducing the blurring effect, allowing the text to be represented more clearly.
@@ 5.1.2 Occupancy Volume Representation
For this task, we maintain the same architectural settings as the image representation task (except we set linear layers ranks as 32 for efficiency) and followed the baseline code of [32] with experimental setup used in FINER [19], while our network translates 3D coordinates (M = 3) into signed distance function (SDF) values (N = 1). In this experiment, five shapes from the public Stanford 3D Scanning Repository dataset [24] are used for evaluation. Table 3 presents a quantitative comparison between the proposed method and four baseline approaches. Across the evaluated shapes, SL2A consistently outperforms the baselines, achieving the best results in all cases. Figure 4 provides comparisons between the proposed method and six baselines showing superior reconstruction quality. Specifically, both low-frequency smooth regions (legs and lower-body curves), and the high-frequency rough parts (dragons face) are well represented by SL2A, providing consistent performance as opposed to other methods. This balanced representation is attributed to our method’s synergistic design, which combines the fusion blocks for capturing low-frequency components with a learnable activation block specialized in high-frequency learning, allowing for a comprehensive spectral representation with the highest IOU values.
Table 3: Quantitative comparisons on representing signed distance field, red indicating the best and blue indicating the second best.
@@ 6 Discussion
In this section, we delve into the impact of various design choices on the model’s overall performance. To this end, we conducted a thorough ablation study, examining the effect of polynomials degree, the distribution of eigen values of neural tangent kernel (NTK) [16], the role of the fusion block, and a qualitative analysis of the activation functions learned by our method. Further details and analysis are provided in the supplementary material.
6.1 Chebyshev Polynomials Degree
Fig. 7 demonstrates the impact of different Chebyshev polynomial degrees ($K$) with and without skip-connections (indicated by an asterisk) in our method. The addition of skip-connections shows a significant improvement, enabling the high-frequency information learned in the initial trainable activation layer to propagate more effectively through subsequent layers. Furthermore, increasing the polynomial degree $K$ generally improves performance across images.
Figure 6: Distribution of eigenvalues of NTK.
Figure 7: Effect of Chebyshev polynomial degree and skip-connections on PSNR of image representation task across different configurrations of our model.
@@ 6.2 Neural Tangent Kernel Perspective
where $\nabla_{\theta}f(x;\theta)$ represents the gradient of the network output with respect to its parameters, and the expectation is taken over the random initialization of $\theta$. The NTK eigenvalues play a crucial role in determining the rate at which different components of the target function are learned. Components corresponding to larger kernel eigenvalues are learned faster, which is important for overcoming spectral bias in neural networks [44, 39]. In Fig. 6, we present the NTK eigenvalue distribution [44] for various models, focusing on the 1D input-output case described in Sec. 4. Increasing $K$ in our method reduces the rate of eigenvalue decay, resulting in higher values that enhance the model’s ability to capture high-frequency components, which is beneficial for faster learning. Removing skip connections (denoted by asterisk) accelerates eigenvalue decay, highlighting their role in preserving spectral properties. As shown in Fig. 6, we observe a hierarchy in decay patterns: ReLU exhibits the most rapid decay, followed by SIREN, which maintains stronger spectral properties, and FINER, which decays more gradually but still faster than our method.

### v4

@@ Abstract
Implicit Neural Representation (INR), leveraging a neural network to transform coordinate input into corresponding attributes, has recently driven significant advances in several vision-related domains. However, the performance of INR is heavily influenced by the choice of the nonlinear activation function used in its multilayer perceptron (MLP) architecture. To date, multiple nonlinearities have been investigated, but current INRs still face limitations in capturing high-frequency components and diverse signal types. We show that these challenges can be alleviated by introducing a novel approach in INR architecture. Specifically, we propose SL2A-INR, a hybrid network that combines a single-layer learnable activation function with an MLP that uses traditional ReLU activations. Our method performs superior across diverse tasks, including image representation, 3D shape reconstruction, and novel view synthesis. Through comprehensive experiments, SL2A-INR sets new benchmarks in accuracy, quality, and robustness for INR. Our code is publicly available at GitHub.
@@ 1 Introduction
Implicit neural representations (INRs), which model continuous functions over discrete data points, have recently garnered significant attention for their ability to effectively represent 2D images, 3D shapes, neural radiance fields, and other complex structures [22, 33, 25, 31]. Unlike traditional discrete grid-based methods, INRs focus on training neural networks, particularly Multilayer Perceptrons (MLPs), that utilize continuous element-wise nonlinear activation functions to map continuous input coordinates to their respective output values, thereby delivering a continuous signal representation [33]. Due to their flexible, compact, and efficient design, INRs have garnered immense interest in solving tasks dealing with complex and high-dimensional data. They have demonstrated versatility across a wide range of applications, including computer graphics [22, 23, 6], computer vision [19, 44], inverse problems [37], and signal representations [41, 33]. In contrast to Convolutional Neural Networks (CNNs), INRs are not constrained by locality biases, thus exhibit exceptional generalization abilities, feature a concise formulation, and display enhanced flexibility. However, their potential is constrained by various obstacles that limit the application of INRs. Conventional neural networks such as ReLU-based MLPs struggle to capture high-frequency details accurately in signals, such as intricate textures in images or complex shapes in 3D reconstruction, due to their inherent bias towards learning simpler, lower-frequency patterns, known as spectral or low-frequency bias [26, 42]. The pioneering efforts in mitigating this issue can be roughly grouped into several main categories. One approach involves increasing the complexity of the input data manifold and explicitly extracting high-frequency features, such as by positional encoding [38, 23]. An alternative major category of methods utilizes complicated activation functions to achieve more precise approximations [33, 10, 30]. Despite promising results, these approaches still exhibit some degree of spectral bias [43], primarily learning faster lower-frequency components of the input. This indicates that there is still room for improvement in addressing spectral bias more comprehensively [5]. Moreover, periodic activation functions are particularly vulnerable to initialization methods, experiencing notable performance declines when not strictly following initialization protocols [28]. Therefore, these approaches have difficulty representing fine details in complex shapes, limiting their flexibility across signal types and inverse problems. To overcome these challenges, we propose a novel approach that significantly boosts the hierarchical representation capabilities of INRs. Our method achieves superior performance in high-fidelity reconstructions across a wide array of tasks, including images and intricate 3D structures. Additionally, it addresses complex problems such as Neural Radiance Fields (NeRF).
@@ 2 Related Works
Implicit Neural Representation Recent studies have demonstrated significant success in utilizing neural networks to represent various types of signals. These representations are applied across a range of applications, including images [28, 30, 33], occupancy volumes [24], view synthesis [22, 2, 11], and audio/video signals [36, 7]. However, the widely-used ReLU activation function has limitations in learning low-frequency information, which can affect the performance of standard MLPs [38]. In this context, the ReLU activation function often fails to accurately represent the signal, resulting in the loss of important high-frequency components and leading to a suboptimal overall representation. To address this issue, several solutions have been proposed in the literature. One approach involves encoding the input data with sinusoids [38], which maps the input coordinates to a higher dimensional space. The output of this positional encoding is then used as input to the MLP. Another line of research focuses on modifying the activation functions of the MLP. For instance, using periodic activations, as in SIREN [33], Gaussian activations [28], wavelet-based activations [30], or variable-periodic activation functions [18] can help capture high-frequency details more effectively. However, these methods also come with their own challenges. For example, the representational capacity of SIREN can be highly dependent on the choice of hyperparameters for sinusoidal functions, such as frequency, and can be sensitive to initialization, necessitating careful design choices to avoid random variations [40].
@@
Learnable Activation Functions As previously noted, the methods for mitigating spectral bias in INRs have introduced new challenges. In particular, networks utilizing custom activation functions, such as periodic activations, are highly sensitive to initialization schemes and hyperparameter choices. Specifically, the inherent periodicity of sine functions can cause oscillations and slow down convergence, creating challenges during the training process [18]. Additionally, despite advancements, some degree of spectral bias remains present in these approaches. Consequently, it is critical to develop an adaptive function to handle nonlinearity and complex frequency distribution effectively. Instead of relying solely on custom-designed activation functions, which may have their own limitations, we propose using learnable activation functions that adapt based on the data itself as part of our network. A learnable activation function is a type of activation function in neural networks that is not fixed but is optimized during training. Unlike traditional activation functions, which are predetermined and static, learnable activation functions have parameters that are adjusted through backpropagation [13, 3]. These functions can be represented by various forms, such as polynomials, splines, or sigmoid linear units [13, 9, 27], and are designed to adapt dynamically to the data and learning process, potentially enhancing the network’s ability to capture complex patterns and improve performance. Specifically, the recent Kolmogorov-Arnold Network (KAN) [17] utilizes B-splines for their activation functions. Capitalizing on this line of research, we transform the input coordinates into a high-dimensional space by projecting them onto multiple adaptive Chebyshev polynomial [35] functions of a fixed degree, departing from the fixed activation functions or positional encodings. These polynomials act as learnable activations, allowing the network to adaptively capture a wide spectrum of frequency information by leveraging the spectral properties of Chebyshev polynomials to efficiently represent both low and high-frequency components of the input signal (as they exhibit progressively higher-frequency oscillatory components in higher degrees). Besides, we incorporate a feature interaction mechanism where the Learnable Activation (LA) Block’s output serves as a dynamic modulator for subsequent layers, enhancing the model’s ability to capture and propagate high-frequency features throughout the network. For simplicity, we use SL2A and SL2A-INR interchangeably throughout the paper to refer to our model.
@@ 3.2.2 Fusion Block
Table 1: Comparison of different methods based on PSNR (in dB) / SSIM (0-1) for 2D color image approximation results for images 00 to 15 and the average. The table is split into two parts for clarity. Higher values indicate better quality, with the best result in red and the second best in blue.
[…]
@@
where $g(x)$ is defined over 2048 discrete points, uniformly sampled from the interval [1,1][-1,1]. In this expression, $R$ represents the rounding function, which introduces discontinuities into the data and increases the difficulty of training. This setup enables us to analyze how the network handles different frequency components, particularly in the presence of spectral bias.
@@ 5 Experimental Results
To evaluate the effectiveness of our proposed method, we conducted several experiments in signal representation tasks such as 2D image representation, 3D shape reconstruction, and novel view synthesis. Our experiments were conducted using PyTorch on an Nvidia RTX 4070 GPU equipped with 12GB of memory.
@@ 5.1.1 Image Representation
[…]
@@
Results We evaluated our method by comparing its performance against five baseline models in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [14]. The baselines included WIRE [30], SIREN [33], FINER [18], ReLU+P.E. [38], and Gauss [28]. For consistency, we trained each model for 500 epochs and reported the best results achieved. All models were implemented with a standard architecture of $L=5$ layers, each containing 256 neurons, except for WIRE, which, following its original design, used a multilayer perceptron (MLP) with 4 layers and 300 neurons per layer. In our approach, for a comparable number of parameters, we used a Learnable Activation (LA) Block with polynomial degrees set to $K=256$, and incorporated linear layers with rank 128 in the Fusion Block followed by ReLU activations. This architecture was designed to balance flexibility and computational efficiency, enabling the model to effectively capture high-frequency details while maintaining efficient computation. The quantitative results for images representation task are shown in Tab. 1. Our model performed best, or at least second-best, in most images, achieving state-of-the-art results. In the few instances where the results were comparable, our approach still demonstrated its effectiveness, achieving competitive performance. Furthermore, our model produced the best average results across all images, highlighting its overall effectiveness. We also provide qualitative results in Fig. 3, which shows a representative sample image tested across all baseline methods and our model. The figure contains magnified regions of the images where text is present. The zoomed-in portions show that our approach outperforms the baselines by reducing the blurring effect, allowing the text to be represented more clearly.
@@ 5.1.2 Occupancy Volume Representation
For this task, we maintain the same architectural settings as in the image representation task, except that we set the ranks of the linear layers to 32 for efficiency. We follow the baseline code of [31] with the experimental setup used in FINER [18], while our network translates 3D coordinates (M = 3) into signed distance function (SDF) values (N = 1). In this experiment, five shapes from the publicly available Stanford 3D Scanning Repository dataset [23] are used for evaluation. LABEL:tab:res_sdf presents a quantitative comparison between the proposed method and four baseline approaches. Across the evaluated shapes, SL2A consistently outperforms the baselines, achieving the best results in all cases. Figure 4 provides comparisons between the proposed method and six baselines showing superior reconstruction quality. Specifically, both low-frequency smooth regions (legs and lower-body curves), and the high-frequency rough parts (dragons face) are well represented by SL2A, providing consistent performance as opposed to other methods. This balanced representation is attributed to our method’s synergistic design, which combines the fusion blocks for capturing low-frequency components with a learnable activation block specialized in high-frequency learning, allowing for a comprehensive spectral representation with the highest IOU values.
Table 2: Quantitative comparisons on representing signed distance field.
@@
5.2 Neural Radiance Fields
Neural Radiance Fields (NeRFs) [22] provide a method for synthesizing novel views of complex 3D scenes by representing them as continuous volumetric functions. These functions are optimized using 2D images and corresponding camera poses. The NeRF architecture employs a Multi-Layer Perceptron (MLP) to implicitly encode the scene. It takes 5D input, comprising a 3D spatial location and a 2D viewing direction, and outputs the associated RGB color and volume density. For training, we adopt the same configurations as in previous work [18] for all baselines and report their corresponding results accordingly. To demonstrate SL2A’s capability in capturing high-frequency details, we employ a reduced training set of 25 images rather than the standard 100 images, consistent with previous studies [18, 30]. LABEL:tab:res_nerf presents a quantitative comparison of our method against various methods on the Blender dataset [22]. SL2A outperforms competing approaches in nearly all scenes and remains highly competitive in the few cases where it does not achieve the best results. This demonstrates SL2A’s ability to deliver consistently strong performance across diverse tasks.
Table 3: Quantitative comparisons on novel view synthesis.
Methods Chair Drums Ficus Hotdog Lego Materials Mic Ship
Gauss 32.68 23.16 26.10 32.17 28.29 26.19 33.59 22.28
SIREN 33.31 24.89 27.26 32.85 29.60 27.13 33.28 22.25
WIRE 29.31 22.22 25.91 30.11 25.76 25.05 32.35 21.15
FINER 33.90 24.90 28.70 33.05 30.04 27.05 33.96 22.47
SL2A 34.70 24.33 28.31 33.83 30.63 28.62 33.88 23.43
@@ 6 Discussion
In this section, we delve into the impact of various design choices on the model’s overall performance. To this end, we conducted a thorough ablation study, examining the effect of polynomials degree, the distribution of eigen values of neural tangent kernel (NTK) [15] and the role of the Fusion block. Further details and analysis are provided in the supplementary material.
6.1 Chebyshev Polynomial Degree and the Role of the Fusion Block
Fig. 6(a) demonstrates the impact of different Chebyshev polynomial degrees ($K$) with and without skip-connections (indicated by an asterisk) in our method. The addition of skip-connections shows a significant improvement, enabling the high-frequency information learned in the initial trainable activation layer to propagate more effectively through subsequent layers. Furthermore, increasing the polynomial degree $K$ generally improves performance across images.
(a) Effect of Chebyshev polynomial degree and skip-connections on PSNR.
(b) Distribution of eigenvalues of NTK.
Figure 6: Visualizations comparing ablation results and kernel eigenvalues.
@@ 6.2 Neural Tangent Kernel Perspective
where $\nabla_{\theta}f(x;\theta)$ represents the gradient of the network output with respect to its parameters, and the expectation is taken over the random initialization of $\theta$. The NTK eigenvalues play a crucial role in determining the rate at which different components of the target function are learned. Components corresponding to larger kernel eigenvalues are learned faster, which is important for overcoming spectral bias in neural networks [43, 38]. In Fig. 6(b), we present the NTK eigenvalue distribution [43] for various models, focusing on the 1D input-output case described in Sec. 4. Increasing $K$ in our method reduces the rate of eigenvalue decay, resulting in higher values that enhance the model’s ability to capture high-frequency components, which is beneficial for faster learning. Removing skip connections (denoted by asterisk) accelerates eigenvalue decay, highlighting their role in preserving spectral properties. As shown in Fig. 6(b), we observe a hierarchy in decay patterns: ReLU exhibits the most rapid decay, followed by SIREN, which maintains stronger spectral properties, and FINER, which decays more gradually but still faster than our method.